# Human-Object Interaction with Vision-Language Model Guided Relative Movement Dynamics

### v1

@@ 3.3 Policy Learning via RMD
$\mathbf{g}_{t}=\left\{\mathbf{s}_{t}^{\text{rmd}},\mathbf{d}_{t},\mathbf{h}_{t},\mathbf{o}_{t}\right\}.$ (7)
Reward via RMD. The task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$ is the weighted sum of the human root target reward, object root target reward, and RMD reward. When $r^{G}$ exceeds the threshold $\theta$, we switch the $\mathcal{G}_{i}$ to the next goal $\mathcal{G}_{i+1}$.
$r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)=\lambda_{RMD}\cdot r_{\text{RMD}}+\lambda_{h}\cdot r_{h}^{\text{tar}}+\lambda_{o}\cdot r_{o}^{\text{tar}}.$ (8)
The human root target reward $r_{h}^{\text{tar}}$ is designed to encourage the human root position $x_{t}^{h_{\text{root}}}$ to approach the target position $x_{t}^{h_{\text{target}}}$.
$r_{h}^{\text{tar}}=\exp\left(-\left\|x_{t}^{h_{\text{root}}}-x_{t}^{h_{\text{target}}}\right\|^{2}\right).$ (9)
The object root target reward $r_{o}^{\text{tar}}$ is designed to encourage the object root position $x_{t}^{o_{\text{root}}}$ to approach the target position $x_{t}^{o_{\text{target}}}$.
$r_{o}^{\text{tar}}=\exp\left(-\left\|x_{t}^{o_{\text{root}}}-x_{t}^{o_{\text{target}}}\right\|^{2}\right).$ (10)
The RMD reward $r_{\text{RMD}}$ is the weighted sum of $r_{RMD}^{k}$, which guides the relative movement dynamics between each human part and object part to faithfully follow the plan. Specifically, the value of $r_{RMD}^{k}$ is determined by the edge weight $w_{ij}$ from $\mathcal{B}$.
$r_{\text{RMD}}=\sum_{k=1}^{15}\lambda_{k}\cdot r_{RMD}^{k}.$ (11)
This reward function guides the relative movement dynamics between human parts and object parts to faithfully follow the plan. Specifically, the behavior of $r_{RMD}^{k}$ is determined by the weight $w_{ij}$ as follows:
  • When $w_{ij}=0$, $r_{RMD}$ encourages the relative velocity between the two particles to approach zero, or the angle between their relative velocity and relative position to be approximately $0.5\pi$.
  • When $w_{ij}=1$, the goal is for a small relative distance between the particles, minimizing the angle between their relative velocity and relative position.
  • When $w_{ij}=2$, the aim is to increase the relative distance between the particles while still minimizing the relative velocity angle.
  • Finally, when $w_{ij}=3$, it indicates an unstable relative movement trend, where the particles may either approach or move away from each other. In such cases, we directly set the reward to its maximum value.
To balance different reward weights $\lambda$ in Eq. 8 and Eq. 11, we utilize adaptive weights proposed by [45]. See Appendix C of the supplementary for additional details.
@@
Table 2: Comparison with Baselines in Single-Task Scenario. * Indicates that these baselines have been re-implemented based on their original papers or publicly available code to ensure a consistent experimental setup. Our method achieves state-of-the-art quantitative results in both static and dynamic interaction tasks.
4 Experiment
Our experiments are organized into two main sections: the first evaluates HOI skill learning in a single-task scenario (Sec. 4.1), and the second evaluates it in a long-horizon multi-task scenario (Sec. 4.2).
Implementation Details.
4.1 Comparisons in a Single-Task Scenario
@@
We select a set of interaction tasks that are commonly used in prior works, as well as some novel tasks that have not been previously reported. These include three static HOI tasks: reaching, sitting, and lying down, and three dynamic HOI tasks: carrying, pushing, and opening. While previous works [9, 45] primarily focus on the approach and interaction with objects, they often overlook a critical aspect: after interacting with an object, the agent must return to a neutral state to facilitate subsequent interactions. For instance, in the sitting task, previous methods consider the task complete once the agent is seated. However, this disregards the need for the agent to return to a standing position in order to transition smoothly to the next task. Therefore, we redefine the task completion criteria by introducing a “leaving” step: after interacting with an object, the agent must stand up and walk to a designated position. This step ensures that the agent returns to a neutral pose, which is essential for smooth transitions between tasks. In line with prior works, we use static, movable, and articulated objects from PartNet [23], 3D-Front [6], CIRCLE [2], SAMP [8], and PartNet-Mobility [44]. For each episode, objects are initialized with a random orientation (ranging from 0 to $2\pi$), a random distance (from 4 to 10 meters), and a random scale (from 0.8 to 1.2). The finish position is then randomly sampled at a point 3 meters away from the object. To ensure a fair comparison with previous state-of-the-art methods, we modify the implementations of AMP [27], InterPhys [9], and UniHSI [45] by adjusting their goal states and corresponding reward functions.
@@
In line with prior works [9, 45], we report the Success Rate to measure the percentage of trials in which the character successfully interacts with the object. A trial is considered successful if the distance between the object’s root and a specific humanoid part is less than 20 cm (for tasks like reaching, sitting, or lying down) or if the distance between the object’s root and the target location is less than 20 cm (for tasks like carrying, pushing, or opening). Additionally, we report the Completion Rate, which measures the percentage of trials in which the character successfully completes the interaction with the object and returns to a neutral standing pose (with the character’s root within 20 cm of the finish position). Furthermore, we utilize Precision to quantify the average distance between a specific human body part and its corresponding target position over the course of the entire interaction in successfully completed trials. All metrics are evaluated over 4096 trials per task.
@@
As shown in Tab. 2, our method achieves higher or comparable performance across all these metrics compared to the baselines. Our method outperforms tasks involving the recovery process of static human-object interactions, such as getting up after sitting or lying down, as the Relative Movement Dynamics guide each body part to move away from the object. In contrast, as shown in Fig. 3, UniHSI [45] struggles to get up because it treats the human-object interaction as a sequence of independent spatial reaching tasks, neglecting the temporal dynamics that coordinate the movements of different body parts. Similarly, InterPhys [9] produces unnatural and unstable motions to complete tasks, such as abrupt kicking or thrusting forward, due to its lack of fine-grained spatial-temporal guidance for coordinating the movements of different body parts. Instead, our design ensures that the agent keeps both hands and arms relatively stationary against the back of the sofa. More qualitative visualization videos can be found in the supplementary materials.
@@
UniHSI[45] 31.2 - - 61.3 - - 10.2 - -
$\text{InterPhys}^{*}$[9] 11.4 47.8 27.5 37.3 61.9 54.1 13.8 18.7 16.9
Ours 75.1 71.2 53.8 86.2 84.3 71.8 7.7 13.0 11.2
RMD-LLM 62.8 53.1 39.9 81.7 78.3 67.2 8.9 15.2 13.8
Table 3: Comparison with Baselines in Long-Horizon Multi-Task Scenarios. * Indicates the use of a vanilla combination and switching of different task policies, as the original InterPhys framework cannot handle multi-task scenarios.
@@
We perform an ablation study on key components to demonstrate the impact of the proposed RMD. In the multi-one setup, we simplify the object representation by removing the division into distinct parts based on geometry and interaction types. Instead, we treat the object as a single entity. This modification leads to slight performance degradation as the model loses a finer-grained understanding of object geometry. Instead, we represent the object’s pose and kinematic information using the root pose and its kinematic data. This leads to a slight degradation in model performance, as it lacks a more fine-grained perception of object geometry. In the one-multi setup, we constrain our RMD Planner to model only the dynamics of a single human part interacting with one of the multiple object parts in each planned substep. This configuration is somewhat similar to UniHSI [45], and it results in an overall decline in performance metrics since setting a target for only a single body part disrupts the overall coordination of human movement and is likely to cause the optimization process to become stuck in a local optimum. The results demonstrate that our multi-multi objective significantly enhances performance in complex and challenging interaction scenarios. Furthermore, removing kinematic relation encoding($\tilde{p}^{ij}_{t}$, $\tilde{v}^{ij}_{t}$) or dynamic encoding($w^{\prime ij}_{t}$) harms performance since these components capture the spatial-temporal relationship between the human and the object during interaction. Overall, the proposed RMD plays an important role in ensuring the humanoid interacts completely with the object: approaching, interacting, and returning to an appropriate initial state.
4.2 Comparisons in a Multi-Task Scenario
While the single-task experiments provide valuable insights into the agent’s performance in handling individual human-object interactions, they focus on isolated tasks. However, real-world applications require the agent to manage multiple tasks either concurrently or sequentially, demanding more sophisticated coordination and planning. To evaluate the robustness of our method in such complex settings, we extend our experiments to a multi-task scenario. In this setup, the agent must handle various tasks within a single plan, which challenges the scalability and adaptability of the proposed Relative Movement Dynamics framework. This extension tests the ability of RMD to manage long-horizon, multi-task interactions. To support training and evaluation in this context, we introduce a new dataset, InterPlay, which contains thousands of interaction plans involving static, dynamic, and articulated objects.
InterPlay Dataset.
We utilize high-quality 3D object assets from existing datasets, including PartNet [23], 3D-FRONT [6], SAMP [8], CIRCLE [2], and PartNet-Mobility [44]. These datasets offer a diverse and detailed collection of 3D models that form the backbone of our experimental framework. We manually verified and further annotated the parts and corresponding point clouds of these 3D assets to ensure data consistency and high fidelity. During both training and inference, we randomly sampled at least two objects for interaction, while optionally including additional objects to provide a more comprehensive scene context. For each object, we randomly assigned its size, initial position, and rotation, thereby generating a wide range of spatial layouts that mimic real-world variability. We then rendered top-view images of the layout context and employed our RMD Planner to generate detailed, context-aware plans based on the given instructions. In total, we generated 1,210 plans, encompassing a wide variety of human-object interactions—including static, dynamic, and articulated objects—with varying temporal horizons within an indoor layout.
Settings.
We split our dataset into three categories: static-interaction, dynamic-interaction, and hybrid setups. In the static-only setup, we sequentially interact with at least two static objects as planned in the dataset. In the dynamic-only setup, we sequentially interact with at least two dynamic objects or articulated objects, also following the pre-planned dataset structure. In the hybrid setup, we select at least three objects, ensuring that at least one dynamic human-object interaction task is included in the overall plan. Since UniHSI [45] supports only sequential interactions with static objects, we adapt their released prompt to generate plans based on scenarios from our InterPlay-static-only dataset. InterPhys [9] supports both dynamic and static interaction tasks, but it is a task-specific policy. For a fair comparison, we combine multiple tasks in a vanilla setup to train a multi-task policy. We use one-hot encoding to represent different tasks and switch to the next goal based on heuristic rules.
Metrics.
Following Sec. 4.1, we report the Completion Rate, which measures the percentage of trials in which the character can sequentially complete all sub-step tasks and return to a natural standing pose, with the character’s root within 20 cm of the target position. We introduce the Sub-step Completion Ratio, which measures the ratio of completed sub-steps to the total number of planned sub-steps. This metric provides a more granular view of the agent’s ability to handle complex tasks with multiple sub-goals. Additionally, we use Sub-step Precision to measure the average distance between a specific human part and its corresponding target position, or between the object root and its corresponding target position.
Result Analysis.
Our method benefits from a unified formulation of interaction and fine-grained spatial-temporal guidance for all human body parts, enabling seamless transitions to a multi-task setup. This design not only streamlines the training process by maintaining consistent representations across various tasks but also significantly improves the model’s ability to generalize across different interaction scenarios. The approach achieves state-of-the-art performance in both individual sub-step execution and long-horizon transitions, reflecting its robustness in handling intricate sequences of actions. Without such a unified formulation, the straightforward combination of different tasks [9] often leads to interference between task-specific features and inefficient training dynamics. In contrast, our integrated method preserves critical spatial and temporal information throughout the pipeline. While UniHSI [45] excels at approaching and interacting with static objects, it struggles with actions such as getting up after interaction, which notably limits its performance in long-horizon scenarios. More qualitative visualization videos can be found in the supplementary materials.
Ablation.
While both VLMs and LLMs have extensive world knowledge, VLMs offer more precise spatial awareness owing to their seamless integration of visual and textual data. To explore this further, we modified the input data by replacing top-view images with richly descriptive paragraphs that detail the objects in the scene, outline their various parts, and specify their relative spatial positions. This modification allowed us to rigorously evaluate the performance of a LLM-based planner under conditions lacking explicit visual cues. The results clearly indicate that the spatial perception capability inherent in the vision model plays a critical role in facilitating effective planning.

### v2

@@ 3.2 Automatic Policy Learning via VLM-Guided RMD
$\mathbf{s}_{t}^{\text{RMD}}=\operatorname{concat}_{(i,j)\in E}\left[\tilde{\mathbf{p}}_{ij},\;\tilde{\mathbf{v}}_{ij},\;\mathbf{w}_{ij}^{\prime}\right]\in\mathbb{R}^{\,|E|\times(3+3+4)}.$ (6)
Both $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ generated by VLM are in the form object(spatial-relationship), e.g., armchair(front). For every referenced object we approximate its geometry by an axis-aligned bounding box with edge lengths $l_{x},l_{y},l_{z}$ and centre $\mathbf{c}_{\mathrm{obj}}$. We map each spatial-relationship token $\delta$ to a displacement vector $\Delta\mathbf{q}(\delta)$ in the object’s local frame. For instance, front $\mapsto(0.7\,l_{x},0,0)$. The absolute targets for the human root and the object root are defined as,
$\mathbf{p}^{h}_{\mathrm{tar}}=\mathbf{c}_{\mathrm{obj}}+\Delta\mathbf{q}(\delta_{h}),\qquad\mathbf{p}^{o}_{\mathrm{tar}}=\mathbf{c}_{\mathrm{obj}}+\Delta\mathbf{q}(\delta_{o}).$ (7)
The destinations proposed by the VLM, originally in the global coordinate system, are transformed into the agent-centric frame to produce the final destination state,
$\mathbf{d}_{t}=\{\mathbf{d}_{t}^{h},\mathbf{d}_{t}^{o}\}.$ (8)
To improve local perception and obstacle avoidance, we attach a fixed-resolution heightmap $\mathbf{h}_{t}\!\in\!\mathbb{R}^{9\times 9}$ that records the elevation of nearby geometry in the root frame [61, 13]. In parallel, we describe object dynamics by the axis-aligned bounding-box vertices $\mathbf{V}_{t}^{\text{box}}\!\in\!\mathbb{R}^{8\times 3}$, the rotation angle $\boldsymbol{\theta}_{t}\!\in\!\mathbb{R}^{3}$, the linear velocity $\mathbf{v}_{t}\!\in\!\mathbb{R}^{3}$, and the angular velocity $\boldsymbol{\omega}_{t}\!\in\!\mathbb{R}^{3}$,
$\mathbf{o}_{t}=\bigl{\{}\mathbf{V}_{t}^{\text{box}},\;\boldsymbol{\theta}_{t},\;\mathbf{v}_{t},\;\boldsymbol{\omega}_{t}\bigr{\}}.$ (9)
By concatenating the RMD state $\mathbf{s}_{t}^{\text{RMD}}$, the destination $\mathbf{d}_{t}$, the heightmap $\mathbf{h}_{t}$, and the object state $\mathbf{o}_{t}$, we construct the complete goal representation,
$\mathbf{g}_{t}=\bigl{\{}\mathbf{s}_{t}^{\text{RMD}},\mathbf{d}_{t},\mathbf{h}_{t},\mathbf{o}_{t}\bigr{\}}.$ (10)
Automatic Reward Design via VLM-Guided RMD.
To guide human-object interaction in alignment with the $i$-th planned interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$ produced by the VLM, we design a composite reward function that promotes three key objectives: (i) guiding the human root trajectory towards the spatial target $\mathcal{T}_{H}$, (ii) guiding the object root trajectory towards $\mathcal{T}_{O}$, and (iii) enforcing the relative motion patterns prescribed by the RMD graph $\mathcal{B}$.
To realize these objectives, we define three corresponding reward terms. The term $r_{d}^{\mathrm{h}}$ encourages the human root position $x_{t}^{\mathrm{h}}$ to approach the designated target $\mathbf{d}_{t}^{h}$, and $r_{d}^{\mathrm{o}}$ encourages the object root position $x_{t}^{\mathrm{o}}$ to move toward its designated target $\mathbf{d}_{t}^{o}$,
$r_{d}^{\mathrm{h}}=\exp\left(-\left\|x_{t}^{\mathrm{h}}-\mathbf{d}_{t}^{h}\right\|^{2}\right),\qquad r_{d}^{\mathrm{o}}=\exp\left(-\left\|x_{t}^{\mathrm{o}}-\mathbf{d}_{t}^{o}\right\|^{2}\right).$ (11)
The term $r_{\text{RMD}}$ encourages each edge $e_{ij}\in\mathcal{B}$, representing a (human-part, object-part) pair, to follow the relative motion dynamics specified by its associated edge weight $w_{ij}$, as planned by the VLM. Specifically, $r_{\text{RMD}}$ is computed by summing the individual alignment reward $r_{\text{rmd}}^{ij}(\cdot)$, each of which measures how well the pair $(p_{h_{i}},p_{o_{j}})$ follows the motion pattern defined by $w_{ij}$, weighted by a coefficient $\lambda_{ij}$:
$r_{\text{RMD}}=\sum_{(i,j)\in E}\lambda_{ij}\cdot r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right).$ (12)
Further implementation details of $r_{\text{rmd}}^{ij}(\cdot)$ are provided in the appendix. Combining these terms, the overall task reward at time $t$ is given by:
$r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)=\lambda_{\text{RMD}}\cdot r_{\text{RMD}}+\lambda_{h}\cdot r_{d}^{\text{h}}+\lambda_{o}\cdot r_{d}^{\text{o}}.$ (13)
Once the task reward $r^{G}$, which is bounded within the range $[0,1]$, exceeds 0.9, the policy transitions to the next planned interaction step $\mathcal{G}_{i+1}$ at the following timestep. To balance the contributions of different reward terms, we adopt adaptive weighting strategies as proposed by [13], which dynamically adjust the value of $\lambda$ in Eq. 12 and Eq. 13. This design enables the current reward to be evaluated in a balanced manner and facilitates accurate assessment of progress within the current stage, thereby ensuring the appropriate timing for transitioning to the next goal.
The style reward $r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$ is computed by a discriminator $D(\mathbf{s}_{t-10:t})$ based on a window of 10 consecutive states. Combining the task reward $r^{G}$ and the style reward $r^{S}$, the total reward at timestep $t$ is defined as,
$r_{t}=\alpha_{\text{task}}\,r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+\alpha_{\text{style}}\,r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right).$ (14)
4 Experiment
We structure our experiments into two main parts: evaluating HOI skill learning in a complex long-horizon multi-task scenario (Sec. 4.1) and in a simple single-task scenario (Sec. 4.2).
InterPlay Dataset.
To evaluate our framework in a long-horizon multi-task scenario, we construct a novel dataset that includes both static and dynamic interaction tasks under diverse scene contexts. We incorporate high-quality 3D object assets from existing datasets, including PartNet [62], 3D-FRONT [63], SAMP [27], CIRCLE [64], and PartNet-Mobility [65]. Each scene contains at least two interactive objects and a minimum of two furniture assets, providing richer contextual complexity for HOI. Further details on dataset construction are provided in the appendix.
Implementation Details.
Table 2: Comparison with baselines in a long-horizon multi-task scenario.
$\text{InterPhys}^{*}$ [15] 21.3 47.8 27.5 37.3 61.9 54.1 13.8 18.7 16.9
$\text{TokenHSI}^{*}$ [17] 25.2 52.5 36.0 48.1 65.7 60.1 13.1 16.6 14.4
UniHSI [13] 37.2 - - 61.3 - - 10.2 - -
Ours 75.1 71.2 53.8 86.2 84.3 71.8 7.7 13.0 11.2
Ours w/ LLM 62.8 53.1 39.9 81.7 78.3 67.2 8.9 15.2 13.8
4.1 Evaluation in a Long-Horizon Multi-Task Scenario
In this setting, the agent must sequentially interact with multiple objects according to a predefined plan, posing significant challenges to the robustness of VLM-guided motion policy learning and the ability to handle long-horizon transitions.
Settings.
We split our dataset into three categories: static-interaction, dynamic-interaction, and hybrid setups. In the static-only setup, we sequentially interact with at least two static objects as planned in the dataset. In the dynamic-only setup, we sequentially interact with at least two dynamic objects or articulated objects, also following the pre-planned dataset structure. In the hybrid setup, we select at least three objects, ensuring that at least one dynamic human-object interaction task is included in the overall plan. UniHSI [13] supports only sequential interactions with static objects. We adapt its released prompt to generate plans based on scenarios from our static-only dataset. Both InterPhys [15] and TokenHSI [17] support interactions with both static and dynamic objects, but they only cover a subset of HOI tasks. To enable multi-task evaluation, we re-implement their methods by training individual skills in parallel across different environments, manually designing goal states and corresponding reward functions for each. A rule-based finite-state machine is employed to switch between tasks during execution. This adaptation requires substantial manual effort.
Metrics.
We report the Completion Rate, which measures the percentage of trials in which the character can sequentially complete all sub-step tasks and return to a natural standing pose, with the character’s root within 20 cm of the target position. We introduce the Sub-step Completion Ratio, which measures the ratio of completed sub-steps to the total number of planned sub-steps. This metric provides a more granular view of the agent’s ability to handle complex tasks with multiple sub-goals. Additionally, we use Sub-step Precision to measure the average distance between a specific human part and its corresponding target, or between the object root and its corresponding target.
Result Analysis.
As shown in Tab. 2, our approach achieves state-of-the-art performance in both individual sub-step execution and long-horizon transitions, reflecting its robustness in handling intricate sequences of actions. Our method benefits from a unified formulation of interaction and fine-grained spatial-temporal guidance for all human body parts, enabling seamless transitions to a multi-task setup. This design not only streamlines the training process by maintaining consistent representations across various tasks but also significantly improves the model’s ability to generalize across different interaction scenarios. In the absence of a unified formulation, approaches that naively combine multiple tasks, such as InterPhys [15] and TokenHSI [17], often encounter difficulties in smoothly switching between tasks. In contrast, our integrated method preserves critical spatial and temporal information throughout the pipeline. While UniHSI [13] excels at approaching and interacting with static objects, it struggles with actions such as getting up after interaction, which notably limits its performance in long-horizon scenarios. More qualitative visualization videos can be found in the supplementary materials.
Ablation.
While both VLMs and LLMs possess extensive world knowledge, VLMs exhibit superior spatial awareness and motion imagination thanks to their integration of visual and textual inputs. To evaluate this, we replaced top-view images with detailed textual descriptions and assessed an LLM-based planner under purely textual conditions. The resulting performance drop highlights the essential role of VLMs in our framework. Pretrained on diverse vision-language data, the VLM excels at generating semantically meaningful and spatially grounded plans—capabilities critical for skill learning in complex, multi-object environments. Without VLM guidance, the system struggles to produce coherent long-horizon behaviors, leading to lower task success and degraded motion quality.
Table 3: Comparison with baselines in a single-task scenario.
@@ 4.1 Evaluation in a Long-Horizon Multi-Task Scenario
4.2 Evaluation in a Single-Task Scenario
@@
We select a set of interaction tasks from our InterPlay dataset, including both tasks commonly used in prior work and several novel tasks that have not been previously explored. These include three static HOI tasks: reaching, sitting, and lying down, and three dynamic HOI tasks: carrying, pushing, and opening. While previous works [15, 13] primarily focus on the approach and interaction with objects, they often overlook a critical aspect: after interacting with an object, the agent must return to a neutral state to facilitate subsequent interactions. For instance, in the sitting task, previous methods consider the task complete once the agent is seated. However, this disregards the need for the agent to return to a standing position in order to transition smoothly to the next task. Therefore, we redefine the task completion criteria by introducing a “leaving” step: after interacting with an object, the agent must stand up and walk to a designated position. This step ensures that the agent returns to a neutral pose, which is essential for smooth transitions between tasks. For each episode, objects are initialized with a random orientation (ranging from 0 to $2\pi$), a random distance (from 4 to 10 meters), and a random scale (from 0.8 to 1.2). The finish position is then randomly sampled at a point 3 meters away from the object. To ensure a fair comparison with previous state-of-the-art methods, we modify the implementations of AMP [54], InterPhys [15], TokenHSI [17], and UniHSI [13] by adjusting their goal states and corresponding reward functions.
@@
We first introduce a new metric, Completion Rate, to evaluate the percentage of trials in which the character not only completes the interaction with the object but also returns to a neutral standing pose, with the root located within 20 cm of the designated finish position. This metric is designed to reflect the agent’s ability to recover from interaction and complete the full motion sequence, offering a more comprehensive evaluation of control quality. In line with prior works [15, 13], we also report the Success Rate, which measures the percentage of trials in which the character successfully interacts with the object. A trial is considered successful if the distance between the object’s root and a specific humanoid part is less than 20 cm (for tasks like reaching, sitting, or lying down), or if the distance between the object’s root and the target location is less than 20 cm (for tasks like carrying, pushing, or opening). Furthermore, we report Precision to quantify the average distance between a specific human body part and its corresponding target position over the course of the entire interaction in successfully completed trials. All metrics are evaluated over 4096 trials.
@@
As shown in Tab. 3, our method achieves higher or comparable performance across all these metrics compared to the baselines. Our method outperforms tasks involving the recovery process of static human-object interactions, such as getting up after sitting or lying down, as the Relative Movement Dynamics guides each body part to move away from the object. In contrast, as shown in Fig. 3, UniHSI [13] struggles to get up because it treats the human-object interaction as a sequence of independent spatial reaching tasks, neglecting the temporal dynamics that coordinate the movements of different body parts. Similarly, InterPhys [15] produces unnatural and unstable motions to complete tasks, such as abrupt kicking or thrusting forward, due to its lack of fine-grained spatial-temporal guidance for coordinating the movements of different body parts. Instead, our design ensures that the agent keeps both hands and arms relatively stationary against the back of the sofa. More qualitative visualization videos can be found in the supplementary materials.
@@
We perform an ablation study on key components to demonstrate the impact of the proposed VLM-guided RMD. In the multi-one setup, we simplify the object representation by removing the division into distinct parts based on geometry and interaction types. Instead, we treat the object as a single entity. This modification leads to slight performance degradation as the model loses a finer-grained understanding of object geometry. Instead, we represent the object’s pose and kinematic information using the root pose and its kinematic data. This leads to a slight degradation in model performance, as it lacks a more fine-grained perception of object geometry. In the one-one setup, we constrain our VLM-guided RMD Planner to model only the dynamics of a single human part interacting with one of the multiple object parts in each planned substep. This configuration results in an overall decline in performance metrics since setting a target for only a single body part disrupts the overall coordination of human movement and is likely to cause the optimization process to become stuck in a local optimum. The results demonstrate that our multi-multi objective significantly enhances performance in complex and challenging interaction scenarios. Furthermore, removing kinematic relation encoding($\tilde{p}^{ij}_{t}$, $\tilde{v}^{ij}_{t}$) or dynamic encoding($w^{\prime ij}_{t}$) harms performance since these components capture the spatial-temporal relationship between the human and the object during interaction. Overall, the proposed VLM-guided RMD plays an important role.